OpenStudy (anonymous):

A cone with height 4 cm and diameter 4 cm contains water that has a depth of h cm, but the water is dripping out a rate of 1/2 cm^3/s. How fast is the depth of the water decreasing when the depth of the water is 2cm?

OpenStudy (anonymous):

If you're going to reply to this question, thank you in advance! But I need to go to sleep for now. I'll come back to it when I wake up.

jhonyy9 (jhonyy9):

opinion ?

jhonyy9 (jhonyy9):

so if the diameter is 4 cm than how many will be the radius of the base ? after you know it you need to calcule the volum of the cone,hope you know formula for V . so from your words i understand that the water has a depth of h what is equal with the height of cone ,yes ?

OpenStudy (anonymous):



jhonyy9 (jhonyy9):

h why equal r so when h=4 and r=2 ?

OpenStudy (anonymous):

yes so r=h/2

jhonyy9 (jhonyy9):

so ok but where is your answer in secoundum what was the question ?

OpenStudy (anonymous):

dont understand your question

OpenStudy (anonymous):

so when r=h/2 dh/dt=-1/2pi cm^3/s

jhonyy9 (jhonyy9):

how can being this minus ?

OpenStudy (anonymous):

the rate is decreasing

jhonyy9 (jhonyy9):

so in the text of exercise is without minus and without pi ,check it please yes ?

OpenStudy (anonymous):

could be

jhonyy9 (jhonyy9):

so and what will be your final answer for this last question ?

OpenStudy (anonymous):

-1/2pi = -1.57

jhonyy9 (jhonyy9):

this is the time in secoundum like the answer on the last question of this exercise ?

OpenStudy (anonymous):

what does secoundum mean?

jhonyy9 (jhonyy9):

this wann to be seconde

jhonyy9 (jhonyy9):

sorry

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

do you mean the units of the answer?

OpenStudy (anonymous):

1.57cm^3/s

jhonyy9 (jhonyy9):

ok i accept it but depend what choises has Eric in mathbook bye

OpenStudy (anonymous):

bye